# Stochastic Processes and Transport Plans

### v4

@@ 9.1 Conditional Wasserstein Distance
Conditional Wasserstein distances appear already in [26, Section 2.4], [32] as well as in the PhD thesis of Gigli [13, Chapter 4], who studied a metric on the so-called ”geometric tangent space” of the Wasserstein-$2$ space. Recently, conditional Wasserstein distances were studied in the realm of machine learning in [4] as well as in functional optimal transport in [21]. We will focus on its application in Bayesian flow matching as presented in our paper [7], where also the proofs of the following theorems and propositions are provided.
@@
Now we define the conditional Wasserstein(-$2$) distance by
@@
As shown in [24, Theorem 3,4] and [4], there is a counterpart of Theorem 3.3 for the conditional Wasserstein distance.
@@
To proof $\rm{i)}$ and $\rm{ii)}$ follow from [7, Lemma 5, Proposition 6]. The first statements of $\rm{iii)}$ follow from Proposition 4.6 and finally the absolutely continuity with respect to $W_{2,\eta}$ is implied by Theorem 9.3. ∎
@@
The counterpart of Proposition 4.4 reads as follows, see [7, Lemma 5]).
@@ 9.2 Almost Conditional Couplings
One drawback of the space $\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ is that we can in general not approximate $\mu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ by an empirical measure, if $\eta$ is not empirical. In other words, an empirical approximation $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(w_{i},x_{i})}$ of $\mu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ with respect to $W_{2}$ is usually not an element of $\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$. Even if we approximate two measures $\mu,\nu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ with respect to $W_{2}$ by empirical measures $\mu_{n},\nu_{n}$ with $\pi^{1}_{\sharp}\mu_{n}=\pi^{1}_{\sharp}\nu_{n}$, it is not necessarily true that there is a sequence $\alpha_{n}\in\Gamma_{\pi^{1}_{\sharp}\mu_{n},o}(\mu_{n},\nu_{n})$ such that $\alpha_{n}$ converges to an $\alpha\in\Gamma_{\eta,o}(\mu,\nu)$ as [7, Example 9] shows.
@@
defines a metric on $\mathcal{P}_{2}(\mathbb{R}^{m}\times\mathbb{R}^{d})$. It turns out that for $0\ll\beta$ the optimal couplings $\alpha$ in this metric are ”close” to fulfilling $\pi^{1,1}\alpha=\Delta\eta$ in the following sense, see [7, Proposition 10].
@@
The distance $W_{\beta}$ successfully addresses the issue of approximating measures by empirical measures as the following proposition from [7, Proposition 12] shows.
@@ 9.3 Bayesian Flow Matching
ii) We can use Proposition 9.1 ii) and solve the optimal transport problem for fixed samples $w$ from $\eta$, to obtain an optimal coupling $\alpha\in\Gamma_{\eta,o}(\mu_{0},\mu_{1})$. In order to ensure that the $\alpha_{w}\in\Gamma_{o}(\mu_{0}^{w},\mu_{1}^{w})$ can be used to build $\alpha\in\Gamma_{\eta,o}(\mu_{0},\mu_{1})$, some measurability conditions must be ensured. For examples where these conditions are fulfilled, see e.g. [14, Corollary 1.2] and [7, Proposition 8].
@@ 9.4 Numerical Examples of Bayesian Flow Matching
We used the dataset Cifar10 [27], which consists of 10 classes of color images of size $32\times 32\times 3$. Here $\eta=\mathcal{P}_{2}(\mathbb{R}^{10})$ is defined by $\eta=\frac{1}{10}\sum_{i=1}^{10}\delta_{e_{i}}$, where $e_{i}$ is the standard basis of $\mathbb{R}^{1}0$. Then samples from $\mu_{1}$ are of the form $(e_{l(x_{i})},x_{i})$, where $l(x_{i})$ is the class of $x_{i}$. The source measure is $\mu_{0}\coloneqq\eta\times\mathcal{N}(0,1)$. For conditional image generation, we used trained flow matching models from [7]. For training with respect to $W_{2,\beta}$ the minibatch OT training algorithm 2 was used for Figure 11[(a),(b)]. For the plan $\alpha\in\Gamma_{\eta}(\mu_{0},\mu_{1})$ in Figure 11[(c)], we applied the coupling from Remark 9.10 and for training Algorithm 1.
Figure 11 shows generated images, where in row $i$ the samples are generated from $(e_{i},z)$ where $10$ different samples $z\sim\mathcal{N}(0,1)$ are drawn. For $\beta=1$, Figure 11[(a)] illustrates that the generation is not class conditional, i.e. in every row there are images of multiple classes. For $\beta=100$ and for the conditional plan from Remark 9.10, we can see in Figure 11 that the generation is class conditional, meaning that every row only contains samples from one class.
@@
The following example can be found in [7, Section 8.2]. Let $X\in\mathbb{R}^{5}$ be a random variable for which the law $P_{X}$ is a Gaussian mixture model with 10 modes with means uniformly chosen in $[-1,1]^{5}$ and standard derivations $0.1$. We consider the Bayesian inverse problem
@@
The noise variable $\Xi$, which is independent from $X$, is distributed as $\mathcal{N}(0,0.1)$. Then by [17, Lemma 11] the posterior distribution $P_{X|Y=y}$ is a Gaussian mixture model which can be analytically computed. This allows us to sample from $P_{X|Y=y}$.
@@ 10.1 Curves of Probability Measures from ODEs
We start with classical results from the theory of ODEs, for existence and regularity see, e.g., [41, Corollary 2.6, Theorem 2.10].
@@
Let $f$ and $g$ be a solution of (187) with initializations $f(t)=x$, and $g(s)=f(s)=\bm{\psi}(s,t,x)$, respectively. By Theorem 10.1, we know that $f=g$ on $[0,1]$ and therefore
@@ 10.3 Computing Gradients with the Adjoint Method
For the special loss (197) the variable $x$ consists of two parts, namely $(x,y)$ and $\psi$ corresponds to $\Psi$ and thus we can evaluate the gradient of the loss function.
@@ 10.4 Numerical Examples of Continuous Normalizing Flows
where we approximate the expectation by an empirical expectation. In order to compute $a_{0}^{\theta}$ we need to solve (210). The only ingredients for solving this equation are the vector field $v_{t}(x,y,\theta)$, the gradients thereof, and the initial conditions. The velocity field $v_{t}$ and its gradients are available since we parametrize it by a neural network. For the first initial condition we need $\psi(1,x,0,\theta)$ which we can obtain by solving $\partial_{t}f(t)=v_{t}(f(t),\theta),f(0)=(x,0)$. For the second initial condition we need $\nabla_{x,y}F$ for $F(x,y)=y-\log p_{1}(x)$. Thus we need to be given a density where the gradient $\nabla_{x}\log p_{1}$ is tractable, which is the case, e.g., for the standard Gaussian distribution. In total we need samples from $p_{X}$ and a density $p_{1}$ for which $\nabla_{x}\log p_{1}$ is tractable. The computation of $a_{0}^{\theta}(x)$ can then be done by solving (210), for which we used the library [10]. Figure 13 shows the trajectories for a vector field obtained via continuous normalizing flow training with three different target distributions. For Figure 13(a) we chose a GMM with $8$ equally weighted modes. The trajectories differ from the ones obtained by flow matching in Figure 7. For Figure 13(b) resp. Figure 13(c) we chose the moons resp. spirals data set from [35].
@@ 11 A Glimpse at Score-Based Diffusion
Usually, we do not have access to $\nabla\log p_{t}$. Fortunately, by [43, Appendix], the loss function can be rewritten, up to a constant, as
@@
Finally, we like to mention that the above (Fokker-Planck) continuity equation (220) arises from a so-called Wasserstein gradient flow of the KL divergence $KL(p,p_{\text{data}})$, i.e., the velocity field $v_{t}$ is in the subdifferential of the above KL function at $p_{t}$, see e.g. [2, Chapter 10.4].
@@
For a family of subsets $B$ of $X$, we call the smallest monotone class containing $B$ the monotone class generated by $B$. The following theorem can be found, e.g., in [6, Theorem 1.9.3)].

### v5

@@ 9.1 Conditional Wasserstein Distance
Conditional Wasserstein distances appear already in [28, Section 2.4], [34] as well as in the PhD thesis of Gigli [15, Chapter 4], who studied a metric on the so-called ”geometric tangent space” of the Wasserstein-222 space. Recently, conditional Wasserstein distances were studied in the realm of machine learning in [6] as well as in functional optimal transport in [23]. We will focus on its application in Bayesian flow matching as presented in our paper [9], where also the proofs of the following theorems and propositions are provided.
@@
Now we define the conditional Wasserstein(-222) distance by
@@
As shown in [26, Theorem 3,4] and [6], there is a counterpart of Theorem 3.3 for the conditional Wasserstein distance.
@@
To proof $\rm{i)}$ and $\rm{ii)}$ follow from [9, Lemma 5, Proposition 6]. The first statements of $\rm{iii)}$ follow from Proposition 4.6 and finally the absolutely continuity with respect to $W_{2,\eta}$ is implied by Theorem 9.3. ∎
@@
The counterpart of Proposition 4.4 reads as follows, see [9, Lemma 5]).
@@ 9.2 Almost Conditional Couplings
One drawback of the space $\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ is that we can in general not approximate $\mu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ by an empirical measure, if $\eta$ is not empirical. In other words, an empirical approximation $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{(w_{i},x_{i})}$ of $\mu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ with respect to $W_{2}$ is usually not an element of $\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$. Even if we approximate two measures $\mu,\nu\in\mathcal{P}_{\eta}(\mathbb{R}^{m}\times\mathbb{R}^{d})$ with respect to $W_{2}$ by empirical measures $\mu_{n},\nu_{n}$ with $\pi^{1}_{\sharp}\mu_{n}=\pi^{1}_{\sharp}\nu_{n}$, it is not necessarily true that there is a sequence $\alpha_{n}\in\Gamma_{\pi^{1}_{\sharp}\mu_{n},o}(\mu_{n},\nu_{n})$ such that $\alpha_{n}$ converges to an $\alpha\in\Gamma_{\eta,o}(\mu,\nu)$ as [9, Example 9] shows.
@@
defines a metric on $\mathcal{P}_{2}(\mathbb{R}^{m}\times\mathbb{R}^{d})$. It turns out that for $0\ll\beta$ the optimal couplings $\alpha$ in this metric are ”close” to fulfilling $\pi^{1,1}\alpha=\Delta\eta$ in the following sense, see [9, Proposition 10].
@@
The distance $W_{\beta}$ successfully addresses the issue of approximating measures by empirical measures as the following proposition from [9, Proposition 12] shows.
@@ 9.3 Bayesian Flow Matching
ii) We can use Proposition 9.1 ii) and solve the optimal transport problem for fixed samples $w$ from $\eta$, to obtain an optimal coupling $\alpha\in\Gamma_{\eta,o}(\mu_{0},\mu_{1})$. In order to ensure that the $\alpha_{w}\in\Gamma_{o}(\mu_{0}^{w},\mu_{1}^{w})$ can be used to build $\alpha\in\Gamma_{\eta,o}(\mu_{0},\mu_{1})$, some measurability conditions must be ensured. For examples where these conditions are fulfilled, see e.g. [16, Corollary 1.2] and [9, Proposition 8].
@@ 9.4 Numerical Examples of Bayesian Flow Matching
We used the dataset Cifar10 [29], which consists of 10 classes of color images of size 32×32×332$\times$ 32$\times$ 332 × 32 × 3. Here $\eta=\mathcal{P}_{2}(\mathbb{R}^{10})$ is defined by $\eta=\frac{1}{10}\sum_{i=1}^{10}\delta_{e_{i}}$, where $e_{i}$ is the standard basis of $\mathbb{R}^{1}0$. Then samples from $\mu_{1}$ are of the form $(e_{l(x_{i})},x_{i})$, where $l(x_{i})$ is the class of $x_{i}$. The source measure is $\mu_{0}\coloneqq\eta\times\mathcal{N}(0,1)$. For conditional image generation, we used trained flow matching models from [9]. For training with respect to $W_{2,\beta}$ the minibatch OT training algorithm 2 was used for Figure 11[(a),(b)]. For the plan $\alpha\in\Gamma_{\eta}(\mu_{0},\mu_{1})$ in Figure 11[(c)], we applied the coupling from Remark 9.10 and for training Algorithm 1.
Figure 11 shows generated images, where in row $i$ the samples are generated from $(e_{i},z)$ where 101010 different samples $z\sim\mathcal{N}(0,1)$ are drawn. For $\beta=1$, Figure 11[(a)] illustrates that the generation is not class conditional, i.e. in every row there are images of multiple classes. For $\beta=100$ and for the conditional plan from Remark 9.10, we can see in Figure 11 that the generation is class conditional, meaning that every row only contains samples from one class.
@@
The following example can be found in [9, Section 8.2]. Let $X\in\mathbb{R}^{5}$ be a random variable for which the law $P_{X}$ is a Gaussian mixture model with 10 modes with means uniformly chosen in $[-1,1]^{5}$ and standard derivations 0.10.10.1. We consider the Bayesian inverse problem
@@
The noise variable $\Xi$, which is independent from $X$, is distributed as $\mathcal{N}(0,0.1)$. Then by [19, Lemma 11] the posterior distribution $P_{X|Y=y}$ is a Gaussian mixture model which can be analytically computed. This allows us to sample from $P_{X|Y=y}$.
@@ 10.1 Curves of Probability Measures from ODEs
We start with classical results from the theory of ODEs, for existence and regularity see, e.g., [43, Corollary 2.6, Theorem 2.10].
@@
Let $f$ and $g$ be a solution of (187) with initializations $f(t)=x$, and $g(s)=f(s)=\bm{\psi}(s,t,x)$, respectively. By Theorem 10.1, we know that $f=g$ on [0,1][0,1][ 0 , 1 ] and therefore
@@ 10.3 Computing Gradients with the Adjoint Method
[…]
@@ 10.4 Numerical Examples of Continuous Normalizing Flows
where we approximate the expectation by an empirical expectation. In order to compute $a_{0}^{\theta}$ we need to solve (210). The only ingredients for solving this equation are the vector field $v_{t}(x,y,\theta)$, the gradients thereof, and the initial conditions. The velocity field $v_{t}$ and its gradients are available since we parametrize it by a neural network. For the first initial condition we need $\psi(1,x,0,\theta)$ which we can obtain by solving $\partial_{t}f(t)=v_{t}(f(t),\theta),f(0)=(x,0)$. For the second initial condition we need $\nabla_{x,y}F$ for $F(x,y)=y-\log p_{1}(x)$. Thus we need to be given a density where the gradient $\nabla_{x}\log p_{1}$ is tractable, which is the case, e.g., for the standard Gaussian distribution. In total we need samples from $p_{X}$ and a density $p_{1}$ for which $\nabla_{x}\log p_{1}$ is tractable. The computation of $a_{0}^{\theta}(x)$ can then be done by solving (210), for which we used the library [12]. Figure 13 shows the trajectories for a vector field obtained via continuous normalizing flow training with three different target distributions. For Figure 13(a) we chose a GMM with 888 equally weighted modes. The trajectories differ from the ones obtained by flow matching in Figure 7. For Figure 13(b) resp. Figure 13(c) we chose the moons resp. spirals data set from [37].
@@ 11 A Glimpse at Score-Based Diffusion
Usually, we do not have access to $\nabla\log p_{t}$. Fortunately, by [45, Appendix], the loss function can be rewritten, up to a constant, as
@@
Finally, we like to mention that the above (Fokker-Planck) continuity equation (220) arises from a so-called Wasserstein gradient flow of the KL divergence $KL(p,p_{\text{data}})$, i.e., the velocity field $v_{t}$ is in the subdifferential of the above KL function at $p_{t}$, see e.g. [4, Chapter 10.4].
@@
For a family of subsets $B$ of $X$, we call the smallest monotone class containing $B$ the monotone class generated by $B$. The following theorem can be found, e.g., in [8, Theorem 1.9.3)].